OpenStudy (anonymous):

how do you factor out x^3+2x^2-16x-32=0?

myininaya (myininaya):

x^2(x+2)-16(x+2)=0 (x+2)(x^2-16)=0 (x+2)(x-4)(x+4)=0

myininaya (myininaya):

x=-2,4,-4

OpenStudy (anonymous):

factor by grouping maybe? lol

OpenStudy (anonymous):

\[x^{3}+2x^{2}-16x-32=0\]\[(x^{3}+2x^{2})-(16x+32)=0\]\[x^{2}(x+2)-16(x+2)=0\]etc...